# Cosmic topology. Part IVa. Classification of manifolds using machine learning: a case study with small toroidal universes

### v1

@@ Abstract
Non-trivial spatial topology of the Universe may give rise to potentially measurable signatures in the cosmic microwave background. We explore different machine learning approaches to classify harmonic-space realizations of the microwave background in the test case of Euclidean $E_{1}$ topology (the 3-torus) with a cubic fundamental domain of a size scale significantly smaller than the diameter of the last scattering surface. Different machine learning approaches are capable of classifying the harmonic-space realizations with accuracy greater than 99% if the topology scale is half of the diameter of the last-scattering surface and orientation of the topology is known. For distinguishing random rotations of these sky realizations from realizations of the covering space, the extreme gradient boosting classifier algorithm performs best with an accuracy of 88%. Slightly lower accuracies of 83% to 87% are obtained with the random forest classifier along with one- and two-dimensional convolutional neural networks. The techniques presented here can also accurately classify non-rotated cubic $E_{1}$ topology realizations with a topology scale slightly larger than the diameter of the last-scattering surface, if provided enough training data. This work identifies the prospects and the main challenges for developing machine learning techniques that are capable of accurately classifying observationally viable topologies.
@@ 1 Introduction
In this work we present a set of machine learning algorithms to distinguish and classify CMB realizations from different topologies in harmonic-space ($a_{\ell m}$ coefficients) and correlation space ($a_{\ell m}a_{\ell^{\prime}m^{\prime}}^{*}$ for each realization). While the problem of distinguishing different topologies using harmonic-space realizations or CMB maps has been considered before (e.g., in Refs. [30, 31]), to the best of our knowledge this is the first application of machine learning techniques to this problem. We start by numerically generating a set of harmonic-space realizations for the cubic $E_{1}$ topology of different sizes, where by size we refer to the length of the fundamental domain in units of the diameter of the LSS. Specifically, we generate cubic $E_{1}$ realizations in four size classes, $L\in\{0.05,0.1,0.5,\infty\}\times L_{\rm LSS}$, where $L_{\rm LSS}$ is the diameter of the LSS and the last class corresponds to the covering space. To classify the realizations, we test the following machine learning algorithms: random forests, extreme gradient boosting classifier (XGBoost), 1-dimensional (1D) convolutional neural networks (CNNs), 2-dimensional (2D) CNNs, and complex 2D CNNs. In each case we obtain the results for the algorithm trained and tested on randomly Wigner-rotated and non-rotated realizations. Additionally, we present a small set of results for harmonic-space realizations with the length scale of $L\gtrsim L_{\rm LSS}$ in order to set preliminary expectations for the effectiveness of machine learning for classifying large universes with non-trivial topologies.
To be clear, $E_{1}$ with a fundamental domain of small size (i.e., $L\lesssim L_{\rm LSS}$) is observationally excluded [17, 18, 19, 20, 21, 22, 23, 15]. However, our objective in this paper is not to extend current observational limits to larger fundamental domains, but to begin developing the machine learning techniques that may ultimately be necessary to do so, and to identify the challenges that must be addressed. To this end, we do not make explicit use of the circles-in-the-sky signature, which does not extend to domains larger than the LSS. We also do not pursue here the standard Bayesian likelihood approach that was also used to obtain earlier constraints.
@@
The layout of the paper is as follows. In Section 2 we discuss the theoretical background, the process of generating the $E_{1}$ topology realizations, and the general features of the resulting datasets. Section 3 outlines the machine learning algorithms we employ. Section 4.1 summarizes the results obtained for realizations with the size scale $L$ significantly smaller than that of the diameter of the LSS. The results for realizations with the size scale $L\gtrsim L_{\rm LSS}$ are discussed in Section 4.2. Future work and implications for observational searches for cosmic topology are discussed in Section 5. Appendix A contains extra information on the training procedures as well as the settings used to train the algorithms. The methodology for generating the realizations with $L\gtrsim L_{\rm LSS}$ is discussed in Appendix B.
@@ 2.3 Features of the $a_{\ell m}$ realization data
Fig. 3 shows a selection of randomly chosen $\mathcal{C}_{\ell m\ell^{\prime}m^{\prime}}$ realizations, while Fig. 4 shows the analytic covariance matrices, as defined in Eq. 2.6, for different sizes of the $E_{1}$ topology. The figures all show the absolute values of correlation normalized by the $\Lambda\text{CDM}$ power spectrum, i.e., by $\left(C_{\ell}^{\Lambda\mathrm{CDM}}C_{\ell^{\prime}}^{\Lambda\mathrm{CDM}}\right)^{-1/2}$. The key standout feature is the checkerboard pattern observed in Fig. 4, which indicates off-diagonal correlations between different multipoles. The relatively orderly and local distribution of features (i.e., nearby in the $C_{\ell m\ell^{\prime}m^{\prime}}$ matrix given the $(\ell,m)$ ordering) hints at this data format being advantageous over some other ways of representing the data. For example, one could consider training an algorithm directly on the temperature fluctuation HEALPix maps for each topology class (for instance by using DeepSphere [43]). However, while for the small topology scales we are considering here (i.e., $L<L_{\rm LSS}$ for $E_{1}$) the correlations might be most clearly represented in real (or pixel) space, for $L>L_{\rm LSS}$ (which is our ultimate target given existing negative circle search results), we expect correlations to be clearest in (3D) Fourier space, and hence, in harmonic space. For cases where matched circles would not be visible, we expect the information to be very widely spread across pixel pairs, and hence, difficult for CNN filters to capture.
@@ 4.1 $E_{1}$ realizations with $L<L_{\rm LSS}$
Table 2: As in Table 1, but for the neural-network-based results. 1D CNN refers to the 1-dimensional convolutional neural network trained directly on $a_{\ell m}$ data, 2D CNN refers to the ResNet-50 architecture trained on 2-dimensional $\mathcal{C}_{\ell m\ell^{\prime}m^{\prime}}$ data, and 2D CVNN refers to the 2-dimensional complex convolutional architecture trained on $\mathcal{C}_{\ell m\ell^{\prime}m^{\prime}}$ data. The accuracy for each individual class is listed in the confusion matrices of Fig. 8.
As before, the confusion matrices for each dataset and each algorithm illustrate the individual topology class performance. Here we find that the algorithms classify the smaller $E_{1}$ topology classes, i.e., $0.05\times L_{\rm LSS}$ and $0.1\times L_{\rm LSS}$, nearly perfectly for all datasets (rotated and unrotated). When trained on randomly rotated realizations, the smaller classes are still classified with near-perfect accuracy, while classifying the larger ones, i.e., $0.5\times L_{\rm LSS}$ and the covering space classes, is more challenging.
@@ 5 Discussion and conclusions
Finally, we note that a spherical harmonic expansion is not the only possible representation of the CMB sky. An alternative representation, the multipole vector formalism [65], has been used extensively in studies of the CMB anomalies [65, 66, 67]. A distinguishing feature of multipole vectors is that rotations of the coordinate system correspond to a rigid rotation of the set of multipole vectors. It is therefore easy to define rotationally invariant quantities, such as dot products of multipole vectors. These features of the multipole vector formalism warrant further investigation using machine learning, which is left for future work.
@@
In summary, we have demonstrated that machine learning techniques are promising for using CMB maps to distinguish classes of topological manifolds, including distinguishing non-trivial topology from the covering space. Our freedom to rotate our coordinate system is the current principle challenge.
@@ A.1 Random forests and extreme gradient boosting classifier
Figure 10: Main components of the ResNet-50 architecture. The convolutional blocks refer to a series of convolutional layers followed by batch normalization and ReLU activation. The final convolutional layer in the convolutional block is generally appended with a shortcut/skip connection. The identity blocks consist of series of convolutional layers with $1\times 1$ and $3\times 3$ kernels followed by batch normalization and ReLU activations. In addition, the identity blocks are appended by shortcut/skip connections and an addition operation, which combines the original input with the learnt features from the convolutional layers in the block. The final block that is appended to the base ResNet-50 architecture (“FC”) contains two Dense layers with 64 and 32 neurons followed by dropout layers with dropout rates of 0.5. Full technical details on the ResNet architecture are provided in Ref. [53]. Figure adapted from Ref. [73].

### v2

@@ Abstract
Non-trivial spatial topology of the Universe may give rise to potentially measurable signatures in the cosmic microwave background. We explore different machine learning approaches to classify harmonic-space realizations of the microwave background in the test case of Euclidean $E_{1}$ topology (the 3-torus) with a cubic fundamental domain of a size scale significantly smaller than the diameter of the last scattering surface. This is the first step toward developing a machine learning approach to classification of cosmic topology and likelihood-free inference of topological parameters. Different machine learning approaches are capable of classifying the harmonic-space realizations with accuracy greater than 99% if the topology scale is half of the diameter of the last-scattering surface and orientation of the topology is known. For distinguishing random rotations of these sky realizations from realizations of the covering space, the extreme gradient boosting classifier algorithm performs best with an accuracy of 88%. Slightly lower accuracies of 83% to 87% are obtained with the random forest classifier along with one- and two-dimensional convolutional neural networks. The techniques presented here can also accurately classify non-rotated cubic $E_{1}$ topology realizations with a topology scale slightly larger than the diameter of the last-scattering surface, if enough training data are provided. While information compressing methods like most machine learning approaches cannot exceed the statistical power of a likelihood-based approach that captures all available information, they potentially offer a computationally cheaper alternative. A principle challenge appears to be accounting for arbitrary orientations of a given topology, although this is also a significant hurdle for likelihood-based approaches.
@@ 1 Introduction
In this work we present a set of machine learning algorithms to distinguish and classify CMB realizations from different topologies in harmonic-space ($a_{\ell m}$ coefficients) and correlation space ($a_{\ell m}a_{\ell^{\prime}m^{\prime}}^{*}$ for each realization). While the problem of distinguishing different topologies using harmonic-space realizations or CMB maps has been considered before (e.g., in Refs. [30, 31]), to the best of our knowledge this is the first application of machine learning techniques to this problem. We start by numerically generating a set of harmonic-space realizations for the cubic $E_{1}$ topology of different sizes, where by size we refer to the length of the fundamental domain in units of the diameter of the LSS. Specifically, we generate cubic $E_{1}$ realizations in four size classes, $L\in\{0.05,0.1,0.5,\infty\}\times L_{\rm LSS}$, where $L_{\rm LSS}$ is the diameter of the LSS and the last class corresponds to the covering space. To classify the realizations, we test the following machine learning algorithms: random forests, extreme gradient boosting classifier (XGBoost), one-dimensional (1D) convolutional neural networks (CNNs), two-dimensional (2D) CNNs, and complex 2D CNNs. In each case we obtain the results for the algorithm trained and tested on randomly Wigner-rotated and non-rotated realizations. Additionally, we present a small set of results for harmonic-space realizations with the length scale of $L\gtrsim L_{\rm LSS}$ in order to set preliminary expectations for the effectiveness of machine learning for classifying large universes with non-trivial topologies.
To be clear, $E_{1}$ with a fundamental domain of small size (i.e., $L\lesssim L_{\rm LSS}$) is observationally excluded [17, 18, 19, 20, 21, 22, 23, 15]. However, our objective in this paper is not to extend current observational limits to larger fundamental domains, but to begin developing the machine learning techniques that may ultimately be necessary to do so, and to identify the challenges that must be addressed. To this end, we do not make explicit use of the circles-in-the-sky signature, which does not extend to domains larger than the LSS. We also note that an analogous problem has been explored using a Bayesian likelihood approach (see, e.g., Ref. [30]), and while we do not explore such methods here, we find it valuable to compare the results obtained in this work with the corresponding results from the likelihood-based approaches. We compare the two different types of methods in Section 5.
@@
The layout of the paper is as follows. In Section 2 we discuss the theoretical background, the process of generating the $E_{1}$ topology realizations, and the general features of the resulting datasets. Section 3 outlines the machine learning algorithms we employ. Section 4.1 summarizes the results obtained for realizations with the size scale $L$ significantly smaller than that of the diameter of the LSS. The results for realizations with the size scale $L\gtrsim L_{\rm LSS}$ are discussed in Section 4.2. Future work, implications for observational topology searches, and a comparison with likelihood-based approaches are discussed in Section 5. Appendix A contains extra information on the training procedures as well as the settings used to train the algorithms. The methodology for generating the realizations with $L\gtrsim L_{\rm LSS}$ is discussed in Appendix B.
@@ 2.3 Features of the $a_{\ell m}$ realization data
Fig. 3 shows a selection of randomly chosen $\mathcal{C}_{\ell m\ell^{\prime}m^{\prime}}$ realizations, while Fig. 4 shows the analytic covariance matrices, as defined in Eq. 2.6, for different sizes of the $E_{1}$ topology. The figures all show the absolute values of correlation normalized by the $\Lambda\text{CDM}$ power spectrum, i.e., by $\left(C_{\ell}^{\Lambda\mathrm{CDM}}C_{\ell^{\prime}}^{\Lambda\mathrm{CDM}}\right)^{-1/2}$. The key standout feature is the checkerboard pattern observed in Fig. 4, which indicates off-diagonal correlations between different multipoles. The relatively orderly and local distribution of features (i.e., nearby in the $C_{\ell m\ell^{\prime}m^{\prime}}$ matrix given the $(\ell,m)$ ordering) hints at this data format being advantageous over some other ways of representing the data. For example, one could consider training an algorithm directly on the temperature fluctuation HEALPix maps for each topology class (for instance by using DeepSphere [43]). While it is true that for the small topology scales that we are considering here (i.e., $L<L_{\rm LSS}$) the correlations might be clearly represented in the real (or pixel) space, extracting this information is non-trivial. It generally requires specific algorithms (e.g., for finding the matched-circle pairs) and a sufficiently large value of $\ell_{\rm max}$ (generally at least $\ell_{\rm max}\approx 250-300$). Extracting this information with machine learning is also complicated by the fact that the correlated pixels generally appear non-locally in the map and are difficult to capture by the convolutional filters. This is especially true in the regime of $L>L_{\rm LSS}$ (which is our ultimate target given the existing negative circle-search results), where we expect correlations to be the clearest in (three-dimensional) Fourier space, and hence, in harmonic space. For these two reasons, we choose to work specifically with harmonic space realizations with the maximum multipole value in the range of $\ell_{\rm max}=50-100$ (and $\mathcal{C}_{\ell m\ell^{\prime}m^{\prime}}$ realizations with $\ell_{\rm max}=20$).
@@ 4.1 $E_{1}$ realizations with $L<L_{\rm LSS}$
Table 2: As in Table 1, but for the neural-network-based results. 1D CNN refers to the one-dimensional convolutional neural network trained directly on $a_{\ell m}$ data, 2D CNN refers to the ResNet-50 architecture trained on two-dimensional $\mathcal{C}_{\ell m\ell^{\prime}m^{\prime}}$ data, and 2D CVNN refers to the two-dimensional complex convolutional architecture trained on $\mathcal{C}_{\ell m\ell^{\prime}m^{\prime}}$ data. The accuracy for each individual class is listed in the confusion matrices of Fig. 8.
As before, the confusion matrices for each dataset and each algorithm illustrate the individual topology class performance, as shown in Fig. 8. Here we find that the algorithms classify the smaller $E_{1}$ topology classes, i.e., $0.05\times L_{\rm LSS}$ and $0.1\times L_{\rm LSS}$, nearly perfectly for all datasets (rotated and unrotated). When trained on randomly rotated realizations, the smaller classes are still classified with near-perfect accuracy, while classifying the larger ones, i.e., $0.5\times L_{\rm LSS}$ and the covering space classes, is more challenging.
@@ 5 Discussion and conclusions
Furthermore, we note that a spherical harmonic expansion is not the only possible representation of the CMB sky. An alternative representation, the multipole vector formalism [65], has been used extensively in studies of the CMB anomalies [65, 66, 67]. A distinguishing feature of multipole vectors is that rotations of the coordinate system correspond to a rigid rotation of the set of multipole vectors. It is therefore easy to define rotationally invariant quantities, such as dot products of multipole vectors. These features of the multipole vector formalism warrant further investigation using machine learning, which is left for future work.
Techniques presented in this work deal specifically with the harmonic realizations of the full sky maps. Looking forward, an interesting question is whether techniques like this could be applied to real observational data. This raises a number of important questions, such as what would be the effects of working with the harmonic realizations of partial or masked sky maps. Presumably, effects such as these would break isotropy and introduce extra harmonic space correlations, which would be challenging to distinguish from the correlations induced by non-trivial topology. Similarly, effects such as foreground residuals and beam asymmetries would have to be accounted for. While studying these systematics is beyond the scope of this work, it is clear that any successful method applied to real CMB data would have to take into account these observational effects. Hence, correctly accounting for such effects is a natural direction for future work.
The problem of distinguishing harmonic space realizations in non-trivial topology from the corresponding covering space realizations has previously been explored using Bayesian and frequentist approaches in Ref. [30]. While we did not attempt to use likelihood-based methods in this work, the techniques described in Ref. [30] offer a valuable point of comparison. Namely, the authors of Ref. [30] have calculated the likelihood function using $10^{4}$ simulated harmonic space realizations of $E_{1}$ topology of different sizes. The likelihood calculations have then been used to distinguish between two cases of special interest: one of an infinite universe (covering space) and the other corresponding to a universe with an $E_{1}$ topology characterized by the relevant correlation matrix. The statistical significance with which the two scenarios can be distinguished has been quantified by the signal-to-noise ratio statistic (Eq. 12 in Ref. [30]). For instance, a universe with $L\approx 0.66\times L_{\rm LSS}$ (or $T[4,4,4]$ in the units of Ref. [30]) can be distinguished from the covering space with $5\sigma-9.3\sigma$ significance depending on the value of $\ell_{\rm max}$ and the details of the normalization. In our case, for instance with the 1D CNN algorithm, we can distinguish the $L=0.5\times L_{\rm LSS}$ class from the covering space with $86.6\%-88\%$ accuracy for the rotated datasets and $99.5\%-99.6\%$ accuracy for the non-rotated datasets. However, it is important to note that the correlation matrices used in Ref. [30] do not include the integrated Sachs-Wolfe effect (ISW) contribution. Properly accounting for the ISW effect should in principle significantly lower the detection power of the outlined likelihood approach, as the authors have pointed out in their section VIII C. In our work, the used correlation matrices do include the ISW and the Doppler shift contributions.
Nonetheless, the results quoted above indicate that likelihood-based methods can generally distinguish $E_{1}$ from the covering space with statistical significance higher than that of the machine learning methods described in this work. However, it is important to point out the limitations of the likelihood-based approaches. Firstly, the obtained results are very sensitive to the relative orientation of a given $\{a_{\ell m}\}$ set and the corresponding correlation matrix used in the likelihood calculation (this is fundamentally the same issue of dealing with rotated realizations that we have described in this work). The likelihood generally needs to be calculated for over $10^{6}$ orientations to obtain the results quoted in Ref. [30]. Furthermore, a full likelihood search would require considering 17 non-trivial topologies, each with up to 6 real parameters specifying the manifold, up to 3 parameters specifying the location of the observer, and up to 3 further parameters specifying the orientation. Machine learning methods offer a computationally cheaper and more efficient alternative: after a training stage, which generally takes several hours, predictions can be generated for thousands of realizations (and over thousands of orientations) in a matter of seconds.
Another important aspect to consider is that in the likelihood-based approaches it is difficult to know what particular features in harmonic space (or pixel space) carry the information that allows distinguishing non-trivial topology from the covering space. In other words, given a CMB map of non-trivial topology, what information specifically distinguishes it from a corresponding covering space map? Machine learning approaches offer a number of techniques to extract this information. As illustrated in our work, the feature importance analysis provides information on which multipoles (or which angular scales in the sky maps) carry the information that is particularly useful for distinguishing that given realization from the covering space realizations. Such information is particularly valuable when considering $a_{\ell m}$ realizations with $L>L_{\rm LSS}$, as in this regime there are no matched-circle pairs, and hence, it is not trivial to deduce how the information that allows distinguishing between different topologies is encoded in pixel space. In future work we hope to combine the outlined feature importance analysis with other methods such as convolutional neural network layer-wise relevance propagation or saliency analysis. Our hope is to also use autoencoders and information-maximizing neural networks (IMNN’s) [68] to extract features related to non-trivial topology and do likelihood-free inference [69].
@@
In summary, we have demonstrated that machine learning techniques are promising for using harmonic space realizations to distinguish between classes of topological manifolds, including distinguishing non-trivial topology from the covering space. Machine learning methods presented in this work also offer a computationally efficient way of extracting features related to non-trivial topology. Nonetheless, our freedom to rotate our coordinate system is the current principal challenge.
@@ A.1 Random forests and extreme gradient boosting classifier
Figure 10: Main components of the ResNet-50 architecture. The convolutional blocks refer to a series of convolutional layers followed by batch normalization and ReLU activation. The final convolutional layer in the convolutional block is generally appended with a shortcut/skip connection. The identity blocks consist of series of convolutional layers with $1\times 1$ and $3\times 3$ kernels followed by batch normalization and ReLU activations. In addition, the identity blocks are appended by shortcut/skip connections and an addition operation, which combines the original input with the learnt features from the convolutional layers in the block. The final block that is appended to the base ResNet-50 architecture (“FC”) contains two Dense layers with 64 and 32 neurons followed by dropout layers with dropout rates of 0.5. Full technical details on the ResNet architecture are provided in Ref. [53]. Reproduced from Ref. [79]. CC BY 4.0. This ResNet50 image has been obtained by the authors from the Wikimedia website where it was made available by Gorlapraveen123 under a CC BY-SA 4.0 license. It is included withing this article on that basis. It is attributed to Gorlapraveen123.